# COINS: SeMantiC Ids Enhanced COld Item RepresentatioN for Click-through Rate Prediction in E-commerce Search

### v3

@@ 1. Introduction
Existing cold-start solutions fall into two paradigms. Generator-based methods (10.1145/3477495.3531897; 10.1145/3690624.3709336; 10.1007/978-981-97-2262-4_9; 10.5555/3295222.3295434) synthesize cold-item representations via warm-item signals: meta-learning (10.1145/3690624.3709336) pre-trains transferable generators for few-sample adaptation, GANs (10.1145/3477495.3531897) align cold embeddings with warm signal distributions, and VAEs (10.1007/978-981-97-2262-4_9) sample from learned warm-data distributions. Knowledge alignment-based methods (10.1145/3696410.3714852; 10.1145/3626772.3657839; 10.1145/3711858; 10.1145/3474085.3475665) bridge content and collaborative signals: contrastive learning optimizes (10.1145/3474085.3475665) content encoders to match warm collaborative representations, while knowledge distillation (10.1145/3696410.3714852) uses content features as a medium to transfer teacher-model (warm-item) knowledge to student models (cold-item).
@@
Figure 1. Types of Knowledge alignment-based methods.:
To alleviate these issues, this paper proposes a novel item representation fusion and enhancement method based on semantic IDs, termed COINS. Specifically, inspired by OneSearch (chen2025onesearchpreliminaryexplorationunified): a recently deployed end-to-end generative framework that has achieved successful application in e-commerce search. We leverage its proposed RQ-OPQ encoding (endowed with rich hierarchical semantic and collaborative signals) to enhance cold-start item representations. First, for the coarser-grained RQ encoding, we design an adaptive transfer mechanism to enable bidirectional transfer between high-frequency collaborative signals embedded in item IDs and dense semantic information within RQ encodings. Second, for the residual vectors of items obtained via RQ hierarchical quantization (i.e., OPQ encodings), which encapsulate the discriminative information of different items, we integrate such discriminative information into item representations through contrastive learning. Ultimately, the resulting item representations not only incorporate abundant content and collaborative information but also accurately capture the uniqueness of individual items, as illustrate in Figure 1 (c). The main contributions are as follows:
@@ 3.2.3. Baselines
SPM_SID (DBLP:conf/recsys/SinghV0KSZHHWTC24): hashing sub fragments of semantic ids sequences to adaptively obtain the semantics of different items.
DAS (DBLP:journals/corr/abs-2508-10584): Through joint training of semantic collaborative models, use contrastive learning to synchronously optimize the quantization and alignment process of semantic ids and item id.
SaviorRec (yao2025saviorrec): A cold start model based on multimodal information and semantic ids collaborative training to enhance item representation.

### v4

@@ 1. Introduction
Existing cold-start solutions fall into two paradigms. Generator-based methods (Chen et al., 2022; Luo et al., 2025; Kong et al., 2024; Vartak et al., 2017) synthesize cold-item representations via warm-item signals: meta-learning (Luo et al., 2025) pre-trains transferable generators for few-sample adaptation, GANs (Chen et al., 2022) align cold embeddings with warm signal distributions, and VAEs (Kong et al., 2024) sample from learned warm-data distributions. Knowledge alignment-based methods (Zhuang et al., 2025; Xu et al., 2024, 2025; Wei et al., 2021) bridge content and collaborative signals: contrastive learning optimizes (Wei et al., 2021) content encoders to match warm collaborative representations, while knowledge distillation (Zhuang et al., 2025) uses content features as a medium to transfer teacher-model (warm-item) knowledge to student models (cold-item).
@@
Figure 1. Types of Knowledge alignment-based methods.
Enjoying the baseball game from the third-base seats. Ichiro Suzuki preparing to bat.
To alleviate these issues, this paper proposes a novel item representation fusion and enhancement method based on semantic IDs, termed COINS. Specifically, inspired by OneSearch (Chen et al., 2025): a recently deployed end-to-end generative framework that has achieved successful application in e-commerce search. We leverage its proposed RQ-OPQ encoding (endowed with rich hierarchical semantic and collaborative signals) to enhance cold-start item representations. First, for the coarser-grained RQ encoding, we design an adaptive transfer mechanism to enable bidirectional transfer between high-frequency collaborative signals embedded in item IDs and dense semantic information within RQ encodings. Second, for the residual vectors of items obtained via RQ hierarchical quantization (i.e., OPQ encodings), which encapsulate the discriminative information of different items, we integrate such discriminative information into item representations through contrastive learning. Ultimately, the resulting item representations not only incorporate abundant content and collaborative information but also accurately capture the uniqueness of individual items, as illustrate in Figure 1 (c). The main contributions are as follows:
@@ 3.2.3. Baselines
SPM_SID (Singh et al., 2024): hashing sub fragments of semantic ids sequences to adaptively obtain the semantics of different items.
DAS (Ye et al., 2025): Through joint training of semantic collaborative models, use contrastive learning to synchronously optimize the quantization and alignment process of semantic ids and item id.
SaviorRec (Yao et al., 2025): A cold start model based on multimodal information and semantic ids collaborative training to enhance item representation.